OpenStudy (anonymous):

After a sailboat had been on the water for 2 h, a change in the wind direction reduced the average speed of the boat by 6 mph. The entire distance sailed was 64 mi. The total time spent sailing was 4 h. How far did the sailboat travel in the first 2 h? _____________mi

OpenStudy (anonymous):

moneybird will u help me please

OpenStudy (anonymous):

Let x be the original speed. 2x + 2 (x+6) = 64 2x + 2x + 12 = 64 4x + 12 = 64 4x = 52 x = 13 13 * 2 = 26 answer is 26 mi

OpenStudy (anonymous):

that was wrong

OpenStudy (anonymous):

Why?